OpenStudy (anonymous):

The area of the circle BELOW is _____. A. 95.5 cm2 B. 725.8 cm2 C. 2903.3 cm2 D. none of the above

OpenStudy (anonymous):



OpenStudy (anonymous):

well u use the formuldo u kno wat that is

OpenStudy (anonymous):

*formula

OpenStudy (anonymous):

A=xr SQUARED

OpenStudy (anonymous):

Area A=3.14r sq

OpenStudy (anonymous):

right ! @Becca2104

OpenStudy (anonymous):

wow im off..!

OpenStudy (anonymous):

basically u square the radius and multiply it by pie 3.14

OpenStudy (anonymous):

^^^^ yes! Answer is D. 13*13=169*3.14=530.66

OpenStudy (anonymous):

D is the right answer but @Becca2104 u used the wrong numbers

OpenStudy (anonymous):

Answer is 725.4 so D is the right choice :)

OpenStudy (anonymous):

@julie001 LOL oops. :)

OpenStudy (anonymous):

None of the above

OpenStudy (anonymous):

@Becca2104 Good Job though

OpenStudy (anonymous):

AREA of the circle = pi r^2 = pi * (13)^2 = 531.14